OpenStudy (anonymous):

Can someone show me the steps to solve these two equations? 1.solve x+4/5=x+5/7 2.solve 5/x-5=6x

OpenStudy (anonymous):

(x+4) (x+5) ----- = ----- 5 7 This?

OpenStudy (anonymous):

yes

OpenStudy (vjordillo):

1.\[\frac{ x+4 }{ 5 } = \frac{ x+5 }{ 7 }\] By cross multiplication, 7(x+4)=5(x+5) By evaluating and simplifying, 7x+28=5x+25 7x-5x=25-28 2x=-3 x=-2/3

OpenStudy (vjordillo):

Sorry. It should be x=-3/2

OpenStudy (anonymous):

multiply the left by 7/7 multiply the right by 5/5 this will give you a common denominator make your that you factor the numerator properly (7(x+4) = 7x+28)

OpenStudy (anonymous):

So if i changed that to decimal form it would be -1.5 right?

OpenStudy (anonymous):

@ehuman so after that I would need to take the parenthesis out right?

OpenStudy (anonymous):

@ehuman I have to graph this equation but I dont have a graphing calculator :(

OpenStudy (anonymous):

can you plot points on a drawn graph?

OpenStudy (anonymous):

 sorry its crappy, idk if i did it right?

OpenStudy (anonymous):

but i did not crunch the numbers, im assuming you did that

OpenStudy (anonymous):

well i got x=30, x \[\neq 5 or 0\]

OpenStudy (anonymous):

@ehuman

OpenStudy (anonymous):

im assuming you have instructions not listed here, your graph looks (0, 30) not (30,0)

OpenStudy (anonymous):

yea sorry i probably made this more difficult then it should be. The rest of the directions say that to find the solution algebraically to verify the graph and if there are values that are not found in the domain.

OpenStudy (anonymous):

ahhh well you said x ≠5or0 so they are excluded from the domain

OpenStudy (anonymous):

But i am also not sure if that is correct.

OpenStudy (anonymous):

im not sure how to answer based on the information given